You have $160 dollars and save $7 per week.Your friend has $210 dollars and saves $5 dollars a week.In how many weeks will they have saved the same?
Answer by nerdybill(7384)

You have $160 dollars and save $7 per week.Your friend has $210 dollars and saves $5 dollars a week.In how many weeks will they have saved the same?
.
Let w = number of weeks of savings
then
You: 160 + 7w
Friend: 210 + 5w
.
Set them equal to each other and solve for w:
160 + 7w = 210 + 5w
160 + 2w = 210
2w = 50
w = 25 weeks